List of figures within the different sections.

• Section 2, Figure 1: Evaluation state and resources of a local volatile thread • Section 2, Figure 2: Implementing a distributed activity by thread migration • Section 5, Figure 4: Implementing a distributed activity by communicating

objects

• Section 5, Figure 6: Emulating thread migration in a higher-order distributed language

• Section 6, Figure 7: Rebinding ubiquitous resources by means of dynamic linking

• Section 7, Figure 8: Object store representation of a persistent and mobile Tycoon thread

• Section 7, Figure 9: Persistent thread states